OpenStudy (anonymous):

What is an equation of a parabola with the given vertex and focus? vertex: (5, 4) ; focus: (8, 4)

OpenStudy (anonymous):

since y coordinates are same so parabola opens along x axis (y-h)^2=4p(x-k) p=8-5=3 where h=4 and k=5

OpenStudy (anonymous):

Wait wait what now ?

OpenStudy (anonymous):

just put the values of p h and k in the equation (y-h)^2=4p(x-k)

OpenStudy (anonymous):

Whats y ?

OpenStudy (anonymous):

and x ?

OpenStudy (anonymous):

they are always present in the equation . (y-4)^2=12(x-5)

OpenStudy (anonymous):

read the following article complete. http://www.purplemath.com/modules/parabola.htm

OpenStudy (anonymous):

Okay so im about to solve that

OpenStudy (anonymous):

Will my answer be a negitive or a positive ?

OpenStudy (anonymous):

(y-4)^2=12(x-5) is the Answer

OpenStudy (anonymous):

That can't be are you sure its not (y-4)^2=12(x+5)

OpenStudy (whpalmer4):

@Kayy_Drizzyy your formula gives a parabola with a vertex at (-5,4) instead of (5,4) as the problem requests.

OpenStudy (anonymous):

yes agree with whpalmer.

OpenStudy (anonymous):

so would my answer me x 1/12 =(y – 4)2 + 5

OpenStudy (whpalmer4):

If you want it in x= form, it would be \[(y-4)^2=12(x-5)\]\[x=\frac{(y-4)^2}{12}+5\]

OpenStudy (anonymous):

Hold on.

OpenStudy (dumbcow):

haha , apparently this is a multiple choice question

OpenStudy (anonymous):

Yes it is and im getting comfused because im trying to solve it

OpenStudy (anonymous):

For some reason were all getiting different answers.

OpenStudy (anonymous):

Help @sami-21

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

i have been summoned :) well as mentioned above the vertex and focus lie horizontal to each other thus it will be "x = something" then plug in the vertex (5,4) to each of the 2 to check which one works

OpenStudy (anonymous):

would my answer be between A & C ?

OpenStudy (dumbcow):

yes

OpenStudy (whpalmer4):

You don't see the equation Sami and I showed you in the answer list?

OpenStudy (anonymous):

I do im going back over you guys sorry for being slow. I'm just tired and trying to complete this assignment.

OpenStudy (anonymous):

@dumbcow is it A

OpenStudy (whpalmer4):

My old math teacher would say "is that an answer, or a prayer?" :-)

OpenStudy (preetha):

Kayy, dont forget to be courteous and give medals.

OpenStudy (whpalmer4):

Have some faith in your abilities. Make mistakes with confidence, that's what I say :-)

OpenStudy (whpalmer4):

Yes, it is A.

OpenStudy (dumbcow):

i concur ... answer is A

OpenStudy (anonymous):

Thanks you guys

OpenStudy (anonymous):

I wish i could give medals to everyone but unfort. i have to choose one